The University of Texas at Austin

ITS Services Status

Current Service Interruptions

7/31/2014
9:51 PM
RESOLVED: Box service is currently down
Service Restored

The Box service for UT Austin has returned to normal.
The Box status page hasn’t yet been updated as of 21:45, but the service is once again functioning.


Box is currently experiencing a technical issue. The initial description of the issue is as follows:
“As of 6:45 pm PDT, Box is down for all services. Customers may see this in their browser: ‘Sorry we can’t access that page.’
“Our engineers are actively working towards a resolution. We will inform you as soon as we have more information on this issue and a projected time to resolution.
“Updates are also available at status.box.com
“Thank you for your patience.”

7/31/2014
3:46 PM
RESOLVED: Emergency Maintenance on UTLogin
Service Restored

Emergency maintenance is complete. System functionality has returned to normal.

We have determined that this issue affected a portion of UTLogin customers. The duration of the issue was approximately from 3:17PM – 3:29PM.

IAM staff is investigating the root cause of the issue.


Our monitoring system indicates that one of the UTLogin servers is experiencing issues. We will be restarting it shortly.

We have not yet determined if this issue is impacting customers.

IAM staff is investigating the issue. 


We are experiencing issues with UTLogin.  System staff is being engaged.

7/29/2014
4:59 PM
On-Going UT-V Storage Issue
Under Investigation

4:58
The UT-V environment has been stable so far today. UT-V staff continue to monitor the issue and prepare for a rollback to ESXi 5.1 if the driver change does not resolve the issue.
Next scheduled update 8/1/2014 at 16:00.


17:00
The final host’s storage controller driver has been changed.
The UT-V environment has been stable so far today. UT-V staff continue to monitor the issue and prepare for a rollback to ESXi 5.1 if the driver change does not resolve the issue.
Next update 7/29/2014 at 17:00


10:01
The UT-V hosts are in the same state as last night. All UT-V hosts have had their storage controller driver changed except for one host in Commodity 2. The final host has only one VM and we will be working with the owner this morning to gracefully migrate the VM so that we can change the driver on the final host.
The Enterprise cluster (with only ITS VMs) did have a period of high storage latency early this morning and a host connection issue that lasted under 1 minute, however, we believe these are different issues at this point and have a case opened with the vendor.
The UT-V team continues to monitor the situation and a new host issue like we have been seeing will trigger a rollback to ESXi 5.1.


7/27/14 11:36 pm
The UT-V team has completed applying storage controller driver changes for all UT-V hosts with one exception in the Commodity 2 cluster. The final host has only one VM running on it that cannot be migrated. It will be addressed tomorrow after reaching out to the VM’s owner.

Next update by 10:00am.


7/27/14 8:38 pm
Vent17 is restored and affected services have been confirmed functional by service owners.

The UT-V team will be doing rolling reboots (transparent) to apply storage controller driver changes in the Commodity 2 and Enterprise clusters to further stabilize the environment. We do not expect any downtime for VMs.


7/27/14 7:51 pm
Vent17 is being restarted. We expect VMs to be back up by 8:15 at the latest.  Service owners should check their VMs for functionality.

Affected services include:
ApplyTx, Austin Disk, Java Mail, MyBenefits, PYPE, UTWeb, Texas Enterprise Directory


7/27/14 7:24 pm
Affected services for Vent12 include:

Apply Texas, DocRepo (Test), JavaMail, UTWeb, MSSQL, Austin Disk


7/27/14 7:16 pm
Vent12 is scheduled to be restarted at 7:15. We expect VMs to be back up by 7:45 at the latest.  Service owners should check their VMs for functionality.

Vent17 is currently scheduled to be restarted at 7:45.


7/27/14 6:15 pm
We are beginning the rollout of the driver change to the other UT-V clusters. The hosts experiencing issues will be restarted after ITS VM administrators have had a chance to attempt graceful VM shutdowns.


7/27/14 5:55 pm
The hosts have now stopped responding in vCenter. It does appear VMs beyond Austin Disk are affected. This cluster contains only ITS VMs.


Two hosts in the UT-V environment appears to be having storage issues with a particular array. ITS Systems staff are investigating.


7/27/14 5:24 pm
Two hosts in the Enterprise cluster (which has not had the HBA driver change applied) are experiencing storage issues. The known impact is limited to Austin Disk Services at this point but may be wider.


July 27, 2014 2:52pm
The HBA drivers have been changed in the Commodity 1 cluster. It continues to operate stably. UT-V staff are monitoring the environment.


July 25, 2014 8:43pm
UT-V staff continue to work with vendor support to resolve this ongoing issue.  The UT-V team is actively working the following plan:

1. Change the default Host Bus Adapter (HBA) driver for the VMware hypervisor in Commodity 1 cluster and reboot all hosts to register the new driver.
2. Observe cluster functionality to determine if HBA driver change has a positive impact on the current issue.
3a. If no further host drop outs in Commodity 1 cluster, change default HBA’s in remaining UT-V clusters.
3b. If host dropouts continue in Commodity 1 cluster, rollback all UT-V clusters to ESXi 5.1, the last stable install base.

Next update Saturday July 26, 2014, 9pm.


July 25, 2014 6:02pm
The UT-V team continues to roll out a configuration change to attempt to stabilize the most problematic cluster.
Next update at 9:00 PM.


July 25, 2014 3:19pm
All VMs affected this afternoon have been powered back on.
UT-V are working to implement a change we hope will mitigate the issue.
Next Update 6:00 PM


July 25, 2014 2:24pm
The host currently affected by the issue will be rebooted at 3:00pm.


July 25, 2014 2:02pm
Another host is experiencing the storage issue. VM admins will be contacted.


July 25, 2014 11:06am
All affected VM’s have been powered up.


July 25, 2014 7:59am
UT-V staff are investigating a current issue with 3 hosts in UT-V.  Some VMs maybe be non-responsive.  UT-V staff are in the process of compiling lists of affected VMs.


UT-V is experiencing ongoing issues with the DDN storage arrays.  VMware and DDN support groups are engaged.  This alert will remain open for the duration of this ongoing issue. 
Next Update EOD July 25, 2014.

7/25/2014
12:23 PM
University Wiki Service intermittant instability
Service Functioning, Temporary Solution in Place

2:10 pm
The University Wikis Service is experiencing intermittent outages. ITS staff continue to monitor and investigate the root cause of the issue but users should expect slow or unresponsive service. 


1:37pm
ITS Staff are investigating.


1:21pm
As of approximately 1:14 PM, ITS Help Desk staff observed that pages on the University Wiki Service were loading slowly (>45 seconds) or not at all.

Past Service Interruptions

Upcoming Maintenance

<<< 10/06/2013 - 10/26/2013 >>>
Sunday Monday Tuesday Wednesday Thursday Friday Saturday
6

12:00 AM - Blackboard Weekly Maintenance

12:00 AM - Mainframe maintenance window

12:00 AM - Austin Exchange Message Service (AEMS) Planned Maintenance

12:00 AM - Austin Disk Services Planned Maintenance Window

7:00 AM - XML Gateway PROD - prep for UTLogin transition

9:00 AM - VM storage migrations to permit SAN OS upgrades

7

7:00 AM - XML Gateway PROD - prep for UTLogin transition

9:00 AM - VM storage migrations to permit SAN OS upgrades

3:30 PM - PyPE OS Security Patching - Test & Qual

4:30 PM - DEFINE Internals/Enterprise Workflow Expedited Non-Disruptive Maintenance

8

7:30 AM - VPN Maintenance

9:00 AM - VM storage migrations to permit SAN OS upgrades

2:00 PM - Planned DPUSER Maintenance in PROD

9

9:00 AM - VM storage migrations to permit SAN OS upgrades

9:00 AM - uTexas Enterprise Directory (TED) routine maintenance

9:30 AM - Javamail, FTPMail Qual and Test host patching

5:00 PM - Blog Service Weekly Maintenance

10

9:00 AM - VM storage migrations to permit SAN OS upgrades

9:00 AM - Shibboleth Maintenance

10:30 AM - Planned Apollo Maintenance in PROD

2:00 PM - uTexas Enterprise Directory (TED) - Expedited Non-Disruptive

11

7:00 AM - Data Center Cleaning Expedited - Non-Disruptive

9:00 AM - VM storage migrations to permit SAN OS upgrades

5:00 PM - AEMS Office 365: Load balancer change for authentication servers

12

6:00 AM - Footprints Weekly Maintenance

9:00 AM - VM storage migrations to permit SAN OS upgrades

13

12:00 AM - Blackboard Weekly Maintenance

12:00 AM - Austin Exchange Message Service (AEMS) Planned Maintenance

12:00 AM - Austin Disk Services Planned Maintenance Window

12:01 AM - SharePoint 2007 Security Update

9:00 AM - VM storage migrations to permit SAN OS upgrades

14

7:15 AM - Mainframe Green Output destination changing to Austin Disk

9:00 AM - VM storage migrations to permit SAN OS upgrades

15

6:00 AM - VoIP System Patch

6:30 AM - PROD uTexas Identity Manager (TIM) Routine Maintenance

9:00 AM - VM storage migrations to permit SAN OS upgrades

4:00 PM - ERP Web Infrastructure Maintenance Window

6:00 PM - Wiki Service Monthly Maintenance

16

9:00 AM - VM storage migrations to permit SAN OS upgrades

9:30 AM - Javamail, FTPMail PROD quarterly OS patching

5:00 PM - Blog Service Weekly Maintenance

5:00 PM - DEFINE Internals/Enterprise Workflow Planned Maintenance

17

9:00 AM - VM storage migrations to permit SAN OS upgrades

1:00 PM - Urchin Planned Maintenance Window

18

9:00 AM - VM storage migrations to permit SAN OS upgrades

10:00 PM - SharePoint 2010 Maintenance

19

6:00 AM - Footprints Weekly Maintenance

7:00 AM - WebSpace planned database maintenance

9:00 AM - VM storage migrations to permit SAN OS upgrades

20

12:00 AM - Blackboard Weekly Maintenance

12:00 AM - Austin Exchange Message Service (AEMS) Planned Maintenance

12:00 AM - Austin Disk Services Planned Maintenance Window

9:00 AM - VM storage migrations to permit SAN OS upgrades

21

9:00 AM - VM storage migrations to permit SAN OS upgrades

6:00 PM - SBC Patching

22

9:00 AM - VM storage migrations to permit SAN OS upgrades

23

9:00 AM - VM storage migrations to permit SAN OS upgrades

5:00 PM - Blog Service Weekly Maintenance

24

9:00 AM - VM storage migrations to permit SAN OS upgrades

25

6:30 AM - Expedited Non-Disruptive PROD uTexas Identity Manager (TIM) Maintenance

9:00 AM - VM storage migrations to permit SAN OS upgrades

3:00 PM - UTBackup Vendor Maintenance

26

7:00 AM - WebSpace planned database maintenance

9:00 AM - VM storage migrations to permit SAN OS upgrades

9:00 AM - UTLogin routine maintenance

9:00 AM - PROD uTexas Identity Manager (TIM) Routine Maintenance


  No maintenance to be scheduled
  Today's date

We Can Help

Get help from an expert:

* ITS Help and Service Desk

* Call us at 512-475-9400

* Submit a help request online

We also have a walk-in service in the first floor lobby of the Flawn Academic Center (FAC). Stop by and let us help you!